Suppose speeds of vehicles on a particular stretch of roadway are normally distributed with mean 36.6 mph and standard deviation 1.7 mph. Find the probability that the mean speed of 20 randomly selected vehicles is between 35 and 40 mph


1
Expert's answer
2022-01-04T12:00:30-0500

Let "X=" the mean speed of 20 randomly selected vehicles: "X\sim N(\mu, \sigma^2/n)."

Given "\mu=36.6mph, \sigma=1.7mph, n=20"


"P(35<X<40)=P(X<40)-P(X\leq 35)"

"=P(Z<\dfrac{40-\mu}{\sigma/\sqrt{n}})-P(Z\leq \dfrac{35-\mu}{\sigma/\sqrt{n}})"

"=P(Z<\dfrac{40-36.6}{1.7/\sqrt{20}})-P(Z\leq \dfrac{35-36.6}{1.7/\sqrt{20}})"

"\approx P(Z<8.944272)-P(Z\leq -4.209069)"

"\approx1-0.000013\approx0.999987"

The probability that the mean speed of 20 randomly selected vehicles is between 35 and 40 mph is "0.999987."